OpenStudy (anonymous):

i really need help :( please help??!! problem 1: Conditions for an experimental chemistry reaction require a temperature of 298 K. The temperature in the lab is 55°F. Which of the following must you do to meet the requirements? (°F= 9/5 (K - 273) + 32) A. increase the room temperature by 22°F B. decrease the room temperature by 12°C C. decrease the room temperature by 33°F D. increase the room temperature by 18 K **Idk :( please help!!

OpenStudy (anonymous):

#2: Jim, Jane, Ann, and Bill measure an object’s length, density, mass, and volume, respectively. Which student’s measurement might be in liters? A. Bill’s B. Jim’s C. Jane’s D. Ann’s **My answer: A. Bill's is this one right?

OpenStudy (aaronq):

1. (°F= 9/5 (K - 273) + 32) K=298 plug into formula.. compare to 55.. you should be able to do this simple arithmetic

OpenStudy (aaronq):

2. you are right, volume is measured in liters, or in a measurement of length cubed. eg. cm^3 which equals 1 mL.

OpenStudy (anonymous):

so #1 is this?? 9/5(298-273)+32=77 so increases by 22 degrees. so my answer to #1 is this?? A??? and #2 yay :)

OpenStudy (anonymous):

@aaronq ? :)

OpenStudy (aaronq):

yep! (y)

OpenStudy (anonymous):

awesome :) thanks!! :D

OpenStudy (aaronq):

no prob !

OpenStudy (anonymous):

awesome :) thanks!! :D